OpenStudy (anonymous):

Solve the system, using substitution. y = 1 - 5x and 5x + y = 8 no solution infinitely many solutions im pretty sure its no solution

OpenStudy (anonymous):

5x+1-5x=8 1=8

OpenStudy (anonymous):

y=1-5x then 5x+y=8 ll be y=8-5x but from first y=1-5x put in second ll be 1-5x=8-5x

OpenStudy (anonymous):

plug in y into the second equation \[5x+ 1-5x = 8\] notice something? the -5x and 5x cancel out. so yes, there'd be no solution because \[1 \ne 8 \] so \[\emptyset \]

OpenStudy (anonymous):

y = 1 - 5x equation (1) 5x + y = 8 equation (2) just plug in equation 1 into equation 2. Since it already solved for y then just plug in directly (aka) substitude 5x + (1 - 5x ) = 8 0x+1=8 1=8 therefore i think no solution.